OpenStudy (anonymous):

What is the slope of the line that passes through (3, -2) and (-1, -1)? -4/1 1/4 4/1 -1/4

OpenStudy (anonymous):

slope=y2-y1/x2-x1 =-1+2/-1-3 =1/-4 =-1/4